SIPB Services List

Write in suggestions or send mail to <arolfe@mit.edu> .

Service  Mailing List  Machine  Maintainers
anxiety-closet  anxiety-maintainers anxiety-closet  mhpower, jhawk
world-wide-weed  world-wide-weed
afs sipb-afsreq ronald-ann marc, warlord, yandros, arolfe
rosebud, reynelda  zacheiss, jweiss, mhbraun, ghudson, kcr
discuss (beacon)  beacon-maintainers  beacon  marthag, mhpower
discuss (charon)  charon-maintainers  charon  jweiss, mhpower
gimp (in graphics) bug-gimp  AFS  yak
gopher  mhpower  charon  mhpower
im locker  bug-im  AFS  aatharuv
IPv6  sipbv6  limekiller  raeburn
lamp  lamp  lamp-www, lamp-www2  keithw
linerva  linerva  vinegar-pot  linerva-root
linux-digests  digests@bed  bedfellow  kolya
macintosh  sipb-bananas  mariel-hemingway
mail-server  owner-mail-server  penguin-lust  arolfe
mbone  jhawk  AFS, portnoy  jhawk
nethack  marduk  turnip-twaddler  yak
news  usenet  bedfellow, beacon  kolya, ghudson, marthag, mhpower
news.answers  news-answers-request penguin-lust
octave locker  ghudson  AFS  ghudson
outland locker  bug-outland  AFS  yandros, bert, shabby, ghudson

jered, jhawk

postscript locker  postscript  AFS  yandros, tlyu, zacheiss
python locker  python  AFS  alex c, glasser
rtfm ftp  rtfm-maintainers  penguin-lust  arolfe
scripts  scripts  old-faithful, better-mousetrap, not-forward, not-backward scripts-root (quentin, andersk, geofft)
sipb dns  sipb-dnsreq  steve-dallas  ingolia
sipb locker  bug-sipb  AFS  jmorzins, jhawk
sipbnet locker  jhawk  AFS  jhawk, elliot, marthag, bjaspan
sipb-voip  sipb-voip  captain-crunch  quentin, geofft
sipb-xen  sipb-xen  black-mesa  sipb-xen
sql  sql  sliced-bread, kitchen-sink  presbrey
stuff (webserver) stuffmaster  dandelion-patch  arolfe, kchen, jtu, zoz, mhpower
tcl locker  bug-tcl  AFS  shabby, nathanw
usenet-addresses  rtfm-maintainers  penguin-lust  ghudson, mhpower
uucp  sipb-uucp  beacon  kolya, warlord, jtkohl
vmware  sipb-vmware  steve-dallas  tibbetts, wjh, kenta, cporter
webster  bug-sipb  charon, tla  jweiss
windows  licks  deathtongue, milo  asedeno
whats  sipb-whats  AFS  aatharuv, jhawk, lgdean, kenta, kchen, krobinso
whichlocker  bug-outland  charon  yak
whois  mhpower  charon  mhpower
xcluster  jweiss, jhawk  lying-naked-in-the-periwinkle  zacheiss, jweiss, jhawk
Backups  sipb-backup  botm, motf  zacheiss, jweiss
3d locker  bug-3d  AFS  nygren
other machines  -maintainers  other  them

       /afs/sipb.mit.edu/admin/text/RCS/service-list.tex,v 1.10 2007/12/17 16:44:29 quentin Exp